Write an equation of the parabola whose vertex is at (-2,3) and whose focus is at (-4,3)
Answer by lwsshak3(11628)

Write an equation of the parabola whose vertex is at (-2,3) and whose focus is at (-4,3).
***
Given parabola opens leftward:
Its basic form of equation: (y-k)^2=-4p(x-h)
axis of symmetry: y=3
p=2 (distance left of vertex on the axis of symmetry)
4p=8
equation: (y-3)^2=-8(x+2)